OpenStudy (anonymous):

integral of -12 / (4+x^2) dx

OpenStudy (anonymous):

I am not sure if this is would apply to a rule or not...but I get confused with these type of problems.

OpenStudy (anonymous):

my brain is malfunctioning and I appear to be useless. I'm very sorry :P @Mashy can you help?

OpenStudy (anonymous):

This appears to be a standard integral.. of the form \[\int\limits_{}^{}\frac{1}{x^2+a^2} dx\] u know the solution to that? or u want me to help u derive it?

OpenStudy (anonymous):

could you help me?

OpenStudy (anonymous):

sure.. use a trignometric substitution.. we ll derive the answer for the general case that i put up.. then u can apply it to your probelm so what trignometric substituition can u think of?

OpenStudy (anonymous):

would it be (1/a)tan-(a/u)+C?

OpenStudy (anonymous):

ohw wow :P.. u directly gave the answer.. did u look it up somewhere? or u just did the whole thing in ur head? :D

OpenStudy (anonymous):

it is a rule that the teacher gave us... so would it be -6xtan-(2/x) +C?

OpenStudy (anonymous):

IT IS (X/2)?

OpenStudy (anonymous):

well yea.. \[\int\limits_{}^{}\frac{1}{x^2+a^2}dx = \frac{\tan^{-1}(\frac{x}{a})}{a}\] so just substitute now :P.. and see for yourself what the answer would be!

OpenStudy (anonymous):

Yes, but then I would have to find a number that I can multiply so it can equal -12 on the top...

OpenStudy (anonymous):

just take -12 out of the integral.. keep it separate and after integrating do the algebra!!..

OpenStudy (anonymous):

and its (X/2)!

OpenStudy (anonymous):

okay! so I get -6tan-(x/2) +C

OpenStudy (anonymous):

yes.. thats right!

OpenStudy (anonymous):

YAY! THANK YOU!

OpenStudy (anonymous):

your welcome!